the conductance of solution measured by two parallel electrodes of area 100 cm squarevseperated by 10 cm was found to be 0.0001 ohm inverse if volume enclosed betwwen two electrode contain 0.1 mole of salt what is the molar conductivity of salt at same concentration

Grade:12

1 Answers

gvs saikiran
4 years ago
conductance = conductivty * cell constant(l/a)
conductivity = 10-4 / 10-1
k=10-3
conductance=(10-3  × 103 )/0.1= 10